# Extreme softening of QCD phase transition under weak acceleration: first principle Monte Carlo results for gluon plasma

### v2

@@ Introduction.
When ions collide, they transfer their kinetic energy to a rapidly expanding plasma fireball, thus experiencing a rapid deceleration  [^1] The deceleration is an acceleration with $a<0$.. The deceleration is caused by the strong longitudinal chromoelectric fields that mediate the interaction between the ions. The strength of these fields is of the order of $E\sim Q_{s}^{2}/g$, where $g$ is the strong coupling constant. The typical deceleration achieved in a collision is argued to be of the order of the gluon saturation scale, $|a|\simeq Q_{s}\simeq 1$ GeV Kharzeev and Tuchin (2005).
@@
We study the effect of a weak acceleration on the gluon plasma in the first-principle approach within the lattice gauge theory. We implement the $T=T(z)$ temperature profile (4) at a fixed $z$-independent spatial volume, noticing that our $t=0$ results can be extended to any $t$ by replacing the $a(z)\to a(t,z)$ in calculated expectation values Becattini and Rindori (2019); Becattini et al. (2021); Selch et al. (2024). The path-integral formalism justifying our setup has been carefully elaborated in Ref. Selch et al. (2024).
@@
The Wilson action (6) with two couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ allows us to fix separately the spacial, ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\sigma}({\beta}_{\sigma},{\beta}_{\tau})$, and temporal, ${\mathfrak{a}}_{\tau}={\mathfrak{a}}_{\tau}({\beta}_{\sigma},{\beta}_{\tau})$, lattice spacings that stand for the physical length of the elementary lattice links in the appropriate directions.  [^2] To avoid confusion, we use the Fraktur font $\mathfrak{a}$ for a lattice spacing and the Italic font $a$ for an acceleration. We use an improved procedure following Ref. Karsch (1982) to simulate the theory on the asymmetric lattices. The technical details of our simulations are described in the Supplemental Material.
We impose periodic boundary conditions along $x,\,y,\,\tau$ directions and keep open boundaries along the acceleration direction $z$. The physical length of the imaginary time circle, $L_{\tau}\equiv N_{\tau}{\mathfrak{a}}_{\tau}(z)=1/T(z)$ is chosen to match the temperature profile (4) corresponding to a constant proper acceleration. In the $z$ direction, temperature varies only on a central segment that occupies half of the volume. We keep two $\delta L_{z}={\mathfrak{a}}_{\sigma}N_{z}/4\sim(1\dots 4)\,{\rm fm}$ thick slices at both sites of the central segment to mitigate the effects of the open boundary conditions along the $z$ direction. It is important to stress that we made the lengths of the transverse spatial directions $L_{i}=N_{i}{\mathfrak{a}}_{\sigma}(z)$ with $i=x,y$ are made $z$-independent by imposing a particular $z$ dependence for the ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ couplings in the lattice action (6) (the details are given in the Supplemental Material).
@@ Results.
with $L_{\rm bulk}={\left\langle|L_{\bm{x}}|\right\rangle}_{\bm{x}}$ averaged over all $\bm{x}$ coordinates.
@@
In Figure 3, we show the susceptibility of the Polyakov loop at various accelerations. To compare this quantity at the non-accelerating case, we introduce the matching susceptibility, which is measured for the homogeneous, $z$-independent gluon matter at temperature $T_{\rm bulk}$ fixed to match $T=T(z,a)$ for chosen $z$ and $a$ of the accelerating gluon matter. For the matching acceleration, we took the moderate value $a\simeq 8$ MeV. Figure 3 shows that a nonzero acceleration makes the phase susceptibility curve (i) wider and (ii) lower, as compared to the matching $a=0$ case, where the bulk variation of temperature is only taken into account. Both properties coherently imply that the accelerating gluon fluid experiences a broad crossover transition to the deconfining phase instead of the first-order phase transition that is expected in the non-accelerating case.
@@
Figure 3: Susceptibility of the Polyakov loop (8) for the accelerating gluon medium (in notations of Fig. 1). The matching $a=0$ curve shows the susceptibility $\chi=\chi(z)$ found at a globally homogeneous temperature $T_{\rm bulk}$ that matches the local temperature $T=T(z,a\simeq 8\,{\rm MeV})$ for each $z$.
@@
To quantify the effect of acceleration on gluon matter, we obtain the critical temperature and the width of the deconfining transition by fitting the Polyakov loop susceptibility by a Gaussian function at each $a$. The fits are shown in Fig. 3 by the solid lines. The obtained phase diagram of the accelerating gluon matter is demonstrated in Fig. 4.
@@
Figure 4: Phase diagram of the hot gluon matter under acceleration in the $(a,T)$ plane. The error bars, given by the vertical lines, are shown together with the widths of the crossovers given by the semi-transparent bands. The inset depicts the density plot of the Polyakov line in the $(x,z)$ plane at a typical gluon configuration, averaged over the $y$ direction.
@@
Despite the physical value of acceleration $a$ being much smaller than the QCD mass scale, $a\ll\Lambda_{\rm QCD}\simeq 100\,{\rm MeV}$, the effect of the acceleration on the phase diagram is unexpectedly substantial. Our results show that even the weakest studied acceleration, $a\simeq 6\,{\rm MeV}$, converts the first-order thermal phase transition in SU(3) Yang-Mills theory into a smooth crossover. As the acceleration increases towards the biggest studied values $a\simeq 27\,{\rm MeV}$, the width of the crossover becomes larger. On the other hand, the acceleration of gluon plasma does not affect the position of the (pseudo)critical temperature of the deconfinement crossover transition, which remains approximately equal, within a few percent accuracy, to the critical temperature of the phase transition the non-accelerating system, $T^{\rm crossover}(a>0)\simeq T^{\rm 1st\,order}_{c}(a=0)$.
@@ Conclusions.
We show that even the weakest acceleration of the order of $a\simeq 6\,{\rm MeV}$ drastically softens the deconfinement phase transition, converting the first-order phase transition of a static hot gluon system to a very soft and broad crossover without noticeable (with an accuracy within a few percent) shifting the position of the corresponding pseudocritical temperature. This unexpected property, which persists up to the maximal studied acceleration $a\simeq 27\,{\rm MeV}$, is correlated coherently with our other observations: (i) the acceleration appears to drive the hot gluon matter residing in the deconfinement phase towards the confining phase; (ii) correspondingly, hot but still confining gluon matter is driven by acceleration closer to the deconfining domain.
The softening effect on the phase transition may have some consequences in the physical environments where (quark) gluon plasma experiences acceleration. One immediate application could be the early Universe. The softening acceleration effects could also hinder the search for the critical line in ongoing experiments in relativistic heavy-ion collisions.
@@ Acknowledgements.
The work of MNC has been supported by the European Union - NextGenerationEU through grant No. 760079/23.05.2023, funded by the Romanian ministry of research, innovation and digitalization through Romania’s National Recovery and Resilience Plan, call no. PNRR-III-C9-2022-I8. VAG and AVM have been supported by RSF (Project No. 23-12-00072, https://rscf.ru/project/23-12-00072/). DVS and ASP were supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia. The numerical simulations were performed at the computing cluster of Far Eastern Federal University and the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS (https://cc.dvo.ru).

### v3

# Extreme Softening of QCD Phase Transition under Weak Acceleration: First-Principles Monte Carlo Results for Gluon Plasma

@@ Introduction.
When ions collide, they transfer their kinetic energy to a rapidly expanding plasma fireball, thus experiencing a rapid deceleration  [^1] The deceleration is an acceleration with $a<0$.. The deceleration is caused by the strong longitudinal chromoelectric fields that mediate the interaction between the ions. The strength of these fields is of the order of $E\sim Q_{s}^{2}/g$, where $Q_{s}{\simeq}1$ GeV is the gluon saturation scale and $g$ is the strong coupling constant. The typical deceleration achieved in a collision is argued to be of the order of the saturation scale, $|a|\simeq Q_{s}$ Kharzeev and Tuchin (2005).
@@
We study the effect of a weak acceleration on the gluon plasma in the first-principle approach within the lattice gauge theory. We implement the $T=T(z)$ temperature profile (4) at a fixed $z$-independent spatial volume, noticing that our $t=0$ results can be extended to any $t$ by replacing the $a(z)\to a(t,z)$ in the calculated expectation values Becattini and Rindori (2019); Becattini et al. (2021); Selch et al. (2024). The path-integral formalism justifying our setup has been carefully elaborated in Ref. Selch et al. (2024).
@@
The Wilson action (6) with two couplings, ${\beta}_{\sigma}$ and ${\beta}_{\tau}$, allows us to fix separately the spatial, ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\sigma}({\beta}_{\sigma},{\beta}_{\tau})$, and temporal, ${\mathfrak{a}}_{\tau}={\mathfrak{a}}_{\tau}({\beta}_{\sigma},{\beta}_{\tau})$, lattice spacings that stand for the physical length of the lattice links in the appropriate directions.  [^2] To avoid confusion, we use the Fraktur font $\mathfrak{a}$ for a lattice spacing and the Italic font $a$ for an acceleration. We use an improved procedure following Ref. Karsch (1982) to simulate the theory on the asymmetric lattices Athenodorou and Teper (2020). The technical details of our simulations are given in the Supplemental Material Chernodub et al. (2025), which includes Refs. Shirogane et al. (2016); Yang et al. (2024).
We impose periodic boundary conditions along $x,\,y,\,\tau$ directions and keep open boundaries along the acceleration direction $z$. The physical length of the imaginary time circle, $L_{\tau}\equiv N_{\tau}{\mathfrak{a}}_{\tau}(z)=1/T(z)$ is chosen to match the temperature profile (4) corresponding to a constant proper acceleration. In the $z$ direction, temperature varies only on a central segment that occupies half of the volume. We keep two $\delta L_{z}={\mathfrak{a}}_{\sigma}N_{z}/4\sim(1\dots 4)\,{\rm fm}$ thick slices at both sites of the central segment to mitigate the effects of the open boundary conditions along the $z$ direction. It is important to stress that we made the lengths of the transverse spatial directions $L_{i}=N_{i}{\mathfrak{a}}_{\sigma}(z)$ with $i=x,y$ are made $z$-independent by imposing a particular $z$ dependence for the ${\beta}_{\sigma}$, and ${\beta}_{\tau}$ couplings in the lattice action (6) (the details are given in Ref. Chernodub et al. (2025)).
@@ Results.
with $L_{\rm bulk}={\left\langle L(z)\right\rangle}_{z}$ computed for the vanishing acceleration $a=0$ and temperature $T_{\rm bulk}$.
@@
In Figure 3, we show the susceptibility of the Polyakov loop at various accelerations. To compare this quantity at the non-accelerating case, we introduce the matching susceptibility, which is measured for the homogeneous, $z$-independent gluon matter at temperature $T_{\rm bulk}$ fixed to match $T=T(z,a)$ for chosen $z$ and $a$ of the accelerating gluon matter. For the matching acceleration, we took the moderate value $a\simeq 5$ MeV. Figure 3 shows that a nonzero acceleration makes the phase susceptibility curve (i) wider and (ii) lower, as compared to the matching $a=0$ case, where the bulk variation of temperature is only taken into account. Both properties coherently imply that the accelerating gluon fluid experiences a broad crossover transition to the deconfining phase instead of the first-order phase transition at zero acceleration.
@@
Figure 3: Susceptibility of the Polyakov loop (8) for the accelerating gluon medium (in the notations of Fig. 1). The matching $a=0$ curve shows the susceptibility $\chi=\chi(z)$ found at a globally homogeneous temperature $T_{\rm bulk}$ that matches the local temperature $T=T(z,a\simeq 5\,{\rm MeV})$ for each $z$.
@@
To quantify the effect of acceleration on hot gluon matter, we fit the Polyakov loop susceptibility with a Gaussian function at each acceleration $a$. We identify the critical temperature with the position of the maximal value of the susceptibility and calculate the corresponding width of the deconfining transition. The fits were obtained by selecting a symmetric region around the maximum selected to saturate the standard deviation $\chi^{2}/{\rm d.o.f}<1$. In Fig. 3, the solid lines represent the best fits for the lattice with the transverse extension $N_{x}=N_{y}=84$. The phase diagram of the accelerating gluon matter is shown in Fig. 4 in the thermodynamic limit (see Chernodub et al. (2025) for details).
@@
Figure 4: Phase diagram of the hot gluon matter under acceleration in the $(a,T)$ plane. The data are presented for the infinite volume limit $N_{x,y}\rightarrow\infty$. The widths of the crossovers are shown by the semi-transparent bands. The inset depicts the density plot of the Polyakov line in the $(x,z)$ plane at a typical gluon configuration, averaged over the $y$ direction.
@@
Despite the physical value of acceleration $a$ being much smaller than the QCD mass scale, $a\ll\Lambda_{\rm QCD}\simeq 100\,{\rm MeV}$, the effect of the acceleration on the phase diagram is unexpectedly substantial. Our results show that even the weakest studied acceleration, $a\simeq 4\,{\rm MeV}$, converts the weak first-order thermal phase transition in SU(3) Yang-Mills theory into a smooth crossover. As the acceleration increases towards the biggest studied values $a\simeq 16\,{\rm MeV}$, the width of the crossover becomes larger. The scaling towards infinite volume, presented in Ref. Chernodub et al. (2025), reaffirms the crossover nature of the transition.
The acceleration of gluon plasma does not affect the position of the (pseudo)critical temperature of the deconfinement crossover transition, which remains approximately equal, within a few percent accuracy, to the critical temperature of the phase transition of the non-accelerating system, $T^{\rm crossover}(a>0)\simeq T^{\rm 1st\,order}_{c}(a=0)$.
@@ Conclusions.
We show that even the weakest acceleration of the order of $a\simeq 4\,{\rm MeV}$ drastically softens the deconfinement phase transition, converting the first-order phase transition of a static hot gluon system to a very soft and broad crossover without noticeably (with an accuracy within a few percent) shifting the position of the corresponding pseudocritical temperature. Our results suggest the existence of a critical acceleration, $a=a_{c}$, at which the weak first-order phase transition of hot SU(3) gluon matter turns into a crossover. The position of the critical point can be probed with the latent heat measurements following, for example, Ref. Shirogane et al. (2016).
The softening effect on the phase transition may have some consequences in the physical environments where (quark) gluon plasma experiences acceleration. One immediate application could be the early Universe. The softening could also hinder the search for the critical line in experiments in relativistic heavy-ion collisions.
@@ Acknowledgements.
The work of MNC was funded by the EU’s NextGenerationEU instrument through the National Recovery and Resilience Plan of Romania - Pillar III-C9-I8, managed by the Ministry of Research, Innovation and Digitization, within the project entitled “Facets of Rotating Quark-Gluon Plasma” (FORQ), contract no. 760079/23.05.2023 code CF 103/15.11.2022. VAG and AVM have been supported by RSF (Project No. 23-12-00072) RSF . AVM thanks the Institut Dennis Poisson (Tours, France) for the kind hospitality and acknowledges the support of Le Studium (Orléans, France) research professorship. DVS and ASP were supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia. The numerical simulations were performed at the computing cluster of Far Eastern Federal University and the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS IAC .